# Emergence and Localisation of Semantic Role Circuits in LLMs

### v1

@@ 5.3 Cross-Scale and Cross-Family Generalisation (RQ3)
We find moderate structural conservation (24–59% node overlap) across scales and model families. Models converge on shared functional vocabularies (high-importance components) while implementing divergent routing patterns: component reuse exceeds connection reuse by $\sim$2:1. Spectral analysis reveals functional alignment despite topological divergence, with small eigenvalue distances (<0.02<0.02) coexisting with 76–88% edge mismatch.
Cross-scale correspondence within Pythia.
@@
⟦t⟧ Model pair Node J. Edge J. $d_{\text{spec}}$ Pythia–14M $\leftrightarrow$ Pythia–410M 0.24 0.12 0.12 Pythia–14M $\leftrightarrow$ Pythia–1B 0.29 0.14 0.11 Pythia–410M $\leftrightarrow$ Pythia–1B 0.44 0.15 0.01 Pythia–1B $\leftrightarrow$ LLaMA–1B 0.59 0.17 0.02
Table 5: Cross-scale and cross-family similarity of role circuits. Node-level overlap increases with model scale, while edge-level overlap remains low, signifying that models tend to reuse similar components while reconfiguring them differently. The cross-family comparison (Pythia–1B $\leftrightarrow$ LLaMA–1B) shows the highest node overlap, suggesting that key components are preserved across architectures even when connectivity patterns diverge. Values report median Top-$K$ Jaccard (K=30K{=}30) and spectral distance ($d_{\text{spec}}$) computed from the lowest k=20k{=}20 Laplacian eigenvalues. Lower $d_{\text{spec}}$ indicates higher functional similarity.
Cross-family correspondence:Pythia-1Bvs. LLaMA-1B.
@@
[…]
6 Discussion
Synthesis across RQs.
Our analyses reveal complementary aspects of semantic-role circuits. RQ1 showed that role information localises into compact, sparsely concentrated subgraphs with role-specific architectural motifs. RQ2 demonstrated that these circuits emerge gradually: indispensability often appears early, but sparsity, density, and attribution patterns evolve smoothly while faithfulness exhibits role-dependent volatility. RQ3 found partial cross-scale and cross-family conservation, with models reusing similar high-importance components but wiring them differently.
Mechanistic interpretation.
Semantic-role circuits appear to arise through prolonged optimisation rather than discrete phase transitions. Although functional competence often emerges early, the underlying computational substrate continues to be reweighted and reorganised. Gradients favour whichever pathways currently yield good predictions, while architectural constraints and implicit regularisation gradually funnel computation into compact motifs (e.g., MLP spines supported by attention bridges). The heterogeneity across roles further indicates that circuits are not simply “switched on” and then fixed: high-importance nodes may stabilise structurally long before the model consistently relies on them. Cross-scale patterns reinforce this view: models share a functional vocabulary, the same components tend to matter—but instantiate it through diverse routing schemes shaped by scale and architectural details.
Linguistic perspective.
From a linguistic standpoint, these circuits approximate role–filler bindings central to the semantics of predicate–argument structure (who did what to whom, when, with what). Their localisation to compact, causally engaged subgraphs suggests that models encode such relations via partially modular mechanisms rather than diffuse token-level heuristics. The partial cross-family convergence, despite different architectures, mirrors a familiar observation in formal semantics: semantic roles behave like typed relational predicates whose abstract relational types remain stable even when different formalisms realise them through distinct derivational routes. By analogy, our results point to shared representational primitives combined with architecture-dependent wiring.
Implications and significance.
Most mechanistic interpretability work investigates narrow behaviours or single checkpoints. Our framework extends this along three axes: (i) temporal, by identifying when circuits first appear and when they become functionally engaged; (ii) cross-scale, by assessing whether mechanisms persist across model sizes and model families; and (iii) semantic, by analysing predicate–argument structure rather than lexical cues or isolated behavioural tasks. The presence of compact, causally functional circuits for our studied semantic structure suggests that LLMs do more than memorise surface co-occurrences. Moreover, partial cross-scale and cross-family transfer indicates that these mechanisms reflect underlying task structure rather than idiosyncratic artifacts, supporting the view that mechanistic interpretability can uncover general computational principles. For safety and alignment, the ability to localise and track abstract semantic mechanisms enables targeted interventions: if specific subgraphs mediate undesirable behaviours, edits can in principle act directly on those circuits rather than global retraining. The heterogeneous emergence timelines we observe also suggest that different semantic capabilities stabilise at different stages, raising the possibility of steering training curricula or early-stopping strategies to emphasise or suppress particular semantic behaviours.
Limitations and future work.
@@ 6 Discussion
Our study focuses on English and a subset of roles; whether similar circuits arise in typologically diverse languages or for richer role inventories remains unknown. The wide change-point confidence intervals reflect the smoothness of structural trajectories rather than methodological noise, but more sensitive emergence metrics (e.g., information-theoretic or curvature-based) may yield finer resolution. Finally, we analyse decoder-only architectures; whether encoder–decoder models exhibit analogous reuse patterns remains an open question. Future work should extend this analysis to multilingual and multimodal models, study interactions between semantic-role circuits and mechanisms for syntax or coreference, and test whether analogous emergence patterns appear for other semantic abstractions such as quantification, modality, or negation.
@@ B.1 Task: Role-Cross Next-Token Prediction
$\begin{split}\text{Accuracy}=\mathds{1}\llbracket y^{(r)}=\operatorname*{arg\,max}_{v}P_{\theta}(v\mid x^{(r)})\rrbracket\\ \land\mathds{1}\llbracket y^{(s)}=\operatorname*{arg\,max}_{v}P_{\theta}(v\mid x^{(s)})\rrbracket\end{split}$ (13)
@@ B.4 Filtering Procedure
We filter generated pairs to retain only examples where the model predicts the role-appropriate target in both contexts.
This ensures circuits are functionally active: the model successfully performs role-specific binding in both clean and corrupted contexts. Examples where either prediction is incorrect are discarded, as they would not reflect active role processing.
@@ B.5 Dataset Statistics
After filtering with Pythia-1B, we obtain:
  • ⟦leftmargin=*, itemsep=1pt⟧
  • Total examples: 4740 role-cross pairs across 8 roles
@@
  • Dual prediction accuracy: 100% (by construction, post-filtering: model predicts correct target in both clean and corrupted contexts)
  • Cross-role coverage: Each role has at least ${\geq}350$ examples
Table 6 provides per-role breakdown.
⟦h⟧ Examples per Role Pythia-14M Pythia-410M Pythia-1B LLaMA-1B Goal 845 895 1052 491 Location 904 673 975 815 Source 667 1206 505 612 Path 802 894 707 959 Instrument 459 1460 1212 1098 Beneficiary 773 837 621 1502 Topic 1120 1323 491 1179 Time 681 645 753 734
@@
Appendix C EAP-IG formalisation
@@
C.1 Transformers as Acyclic Graphs
@@ C.2 EAP-IG Scoring Procedure
    Interpolate inputs: Define the straight-line interpolation $\mathbf{x}(\alpha)=\mathbf{x}^{(r)}+\alpha(\mathbf{x}^{(c)}-\mathbf{x}^{(r)})$ for $\alpha\in\llbracket 0,1\rrbracket$.
@@
Intuitively, $S^{\mathrm{IG}}_{u\to v}$ approximates the first-order change in the loss if the contribution of $u$ to $v$’s input were replaced by its corrupted counterpart, averaged along the naturalistic path from corrupt to clean inputs. By integrating over $\alpha$, the estimate reduces sensitivity to local saturation at the clean point and captures non-linear response accumulated along the path.
C.3 Score Normalisation and Aggregation.
@@
C.4 Circuits Evaluation metrics
@@
Bootstrap CI for Change-point We estimate a split point $t$ in $(x\!=\!\text{step},\,y\!=\!\text{faithfulness})$ by minimising the summed residuals of two OLS lines on $\llbracket x_{1},\dots,x_{t}\rrbracket$ and $\llbracket x_{t+1},\dots,x_{n}\rrbracket$ with a minimum segment length of 3. For a 95% CI, we resample pairs $(x,y)$ with replacement, sort by $x$, re-fit the split, map the bootstrap split $x_{t}^{\star}$ back to the original grid, and take the 2.5/97.5 percentiles over 1,000 replicates. We report the point estimate $\hat{t}_{c}$ and the percentile CI over steps.
Appendix D Computation and parameters specifications
D.1 Hyperparameter Selection
Attribution (EAP-IG).
  • ⟦leftmargin=*, itemsep=2pt⟧
Sparsity and Localization (RQ1).
  • ⟦leftmargin=*, itemsep=2pt⟧
Emergence Dynamics (RQ2).
  • ⟦leftmargin=*, itemsep=2pt⟧
Cross-Scale Comparison (RQ3).
  • ⟦leftmargin=*, itemsep=2pt⟧
Data Filtering.
  • ⟦leftmargin=*, itemsep=2pt⟧
Computational Constraints.
Software Specification.
@@ Appendix E Metric Definitions and Interpretation
⟦Top-$K$ node-mass proportion.⟧
@@
Range $\llbracket 0,1\rrbracket$; higher $\Rightarrow$ stronger sparsity. We report $K\in\{5,10,20\}$.
⟦Top-$P$ coverage.⟧ Minimal $K$ such that $\mathrm{TopK}(K)\geq P$, for $P\in\{0.80,0.90,0.95\}$. Lower $K$ indicates higher sparsity.
⟦Gini coefficient (node mass).⟧ Standard Gini on nonnegative masses; range $\llbracket 0,1\rrbracket$ ($1=$ all mass on one node). Primary comparator for sparse localisation (RQ1).
@@
⟦t⟧ Metric Range High Low Top-$K$ mass $\llbracket 0,1\rrbracket$ concentrated circuit (RQ1) diffuse attribution Top-$P$ coverage ($K$) $\mathbb{N}$ few nodes capture $P$ (sparse) many nodes needed Gini (mass) $\llbracket 0,1\rrbracket$ strong sparsity (RQ1) uniform mass Density $\llbracket 0,1\rrbracket$ saturated links (post-$\hat{t}_{c}$) sparse links Reciprocity $\llbracket 0,1\rrbracket$ feedback motifs feed-forward routing Avg out-degree $\mathbb{R}_{\geq 0}$ broad fan-out narrow fan-out Avg weighted out-degree $\mathbb{R}_{\geq 0}$ strong influence spread weak influence Edge-type mix simplex routing vs. residual balance Bridges $\mathbb{N}$ bottlenecks (ablation targets) redundancy Layer span $\mathbb{N}$ deeper integration shallow circuit Avg betweenness $\llbracket 0,1\rrbracket$ (norm.) coordinator hubs flat routing Top-$K$ Jaccard (step) $\llbracket 0,1\rrbracket$ persistent circuit unstable set Cross-model Jaccard $\llbracket 0,1\rrbracket$ architectural consistency family/scale drift Spectral distance $d_{\text{spec}}$ $\mathbb{R}_{\geq 0}$ similar flow geometry divergent geometry
Table 7: Interpretation guide for graph metrics.
@@
⟦Nodes, Edges.⟧ $|V|$ and $|E|$ of the in-circuit graph.
⟦Density.⟧ We use NetworkX’s directed density:
$\begin{split}\mathrm{density}(G)=\frac{|E|}{|V|(|V|-1)}\in\llbracket 0,1\rrbracket,\end{split}$ (21)
@@
⟦Reciprocity.⟧ Fraction of directed edges participating in reciprocated pairs:
@@
⟦Average out-degree / weighted out-degree.⟧
@@
⟦Edge-type fractions.⟧
@@
⟦Bridges (undirected projection).⟧ Count edges whose removal disconnects the undirected projection of $G$ (structural bottlenecks).
⟦Layer span.⟧
@@
⟦Average betweenness centrality.⟧ $\overline{C_{B}}=\tfrac{1}{|V|}\sum_{v}C_{B}(v)$ on the directed graph (normalized; may use sampling for efficiency).
@@
⟦Detectability $t_{\text{det}}$ (optional).⟧ First step where faithfulness exceeds an early-phase baseline plus $2\sigma$ (baseline/variance from the first $2$ checkpoints), persisting for $\geq 2$ checkpoints.
⟦Indispensability $t_{\text{ind}}$.⟧ Earliest step where ablating the discovered circuit yields a statistically significant performance drop that persists for at least $2$ subsequent checkpoints.
⟦Change-point $\hat{t}_{c}$ (with bootstrap CI).⟧ Two-segment least-squares on $y_{t}\in\{\text{faithfulness},\mathrm{TopK}(K)\}$ with a minimum segment length of $3$. We report $\hat{t}_{c}$ with a nonparametric bootstrap $95\%$ CI.
@@
⟦Node/edge overlap across models.⟧ Mean Jaccard of Top-$K$ node/edge sets between two models, averaged over common checkpoints. (Defaults: K=30K{=}30.)
⟦Spectral similarity.⟧ For symmetrised, weighted Laplacians ($w_{uv}=|s(u\!\leftrightarrow\!v)|$),
@@
⟦Within-model role overlap.⟧ Jaccard of Top-$K$ nodes between roles at fixed steps (specialisation vs. shared scaffolding).
@@
⟦1⟧
Model checkpoints $\{\theta_{t}\}_{t=0}^{T}$, role-cross dataset $D$, role $r\in\mathcal{R}$, top-$K$ threshold
Circuit $\mathcal{C}_{t}^{(r)}$ for each $t$; emergence times $(t_{\text{ind}},\hat{t}_{c},t_{\text{cons}})$
Phase 1: Causal Localisation (EAP-IG)
for each checkpoint $t=0,\ldots,T$ do
  Compute CNP scores $\{\Delta_{\theta_{t}}(y;r,s)\}$ for all $(x^{(r)},x^{(s)})\in D$
  Run EAP-IG to obtain edge attributions $\{S_{u\to v}^{\text{IG}}\}_{e\in\mathcal{E}}$ (Appendix C)
  Normalise: $\tilde{S}_{u\to v}^{\text{IG}}\leftarrow S_{u\to v}^{\text{IG}}/\sum_{e}|S_{e}^{\text{IG}}|$
  Extract circuit: $\mathcal{C}_{t}^{(r)}\leftarrow\text{TopK}(\{|\tilde{S}_{u\to v}^{\text{IG}}|\},K)$
end for
Phase 2: Temporal Monitoring
for each checkpoint $t=0,\ldots,T$ do
  Compute faithfulness $F_{t}$ via ablation (Eq. 6)
  Compute stability $S_{t}$ via Jaccard (Eq. 7)
  Compute structural metrics: Top-$K$ node mass, Gini coefficient
end for
Phase 3: Emergence Detection
Detect indispensability: $t_{\text{ind}}\leftarrow\min\{t:M_{t}(\mathcal{E})-M_{t}(\mathcal{C}_{t})>\epsilon\text{ for }\geq 2\text{ steps}\}$
Estimate functional transition: $\hat{t}_{c}\leftarrow\operatorname*{arg\,max}_{t}R^{2}(\text{PiecewiseLinear}(\{F_{t}\}))$ with bootstrap CIs
Detect consolidation: $t_{\text{cons}}\leftarrow\min\{t:S_{t}\geq 0.6\text{ for }\geq 2\text{ steps}\}$
return $\{\mathcal{C}_{t}^{(r)}\}_{t=0}^{T}$, $(t_{\text{ind}},\hat{t}_{c},t_{\text{cons}})$

### v2

@@ 5.3 Cross-Scale and Cross-Family Generalisation (RQ3)
We find moderate structural conservation (24–51% node overlap) across models. They converge on shared functional vocabularies while implementing divergent routing patterns: component reuse exceeds connection reuse by $\sim$2:1. Spectral analysis reveals functional alignment despite topological divergence, with small eigenvalue distances (<0.02<0.02) coexisting with 76–88% edge mismatch.
Cross-scale correspondence.
@@
Model pair Node J. Edge J. $d_{\text{spec}}$
Pythia–14M $\leftrightarrow$ Pythia–410M 0.24 0.11 0.12
Pythia–14M $\leftrightarrow$ Pythia–1B 0.29 0.12 0.10
Pythia–410M $\leftrightarrow$ Pythia–1B 0.42 0.17 0.01
Pythia–1B $\leftrightarrow$ LLaMA–1B 0.51 0.14 0.02
Table 4: Cross-scale and -family similarity of role circuits. Node-level overlap increases with model scale while edge-level overlap remains low. Cross-family comparison shows the highest node overlap. Values report median Top-$K$ Jaccard (K=30K{=}30) and spectral distance. Lower $d_{\text{spec}}$ indicates higher functional similarity.
Cross-family correspondence.
@@
Figure 5: Cross-family correspondence for Location (top) and Instrument (bottom). Node sets align substantially more than edges between Pythia–1B and LLaMA–1B, suggesting shared component selection but model-specific routing.
@@
(i) Moderate structural conservation across scales and families. High-importance nodes show 24–51% overlap; cross-family correspondence (51%) exceeds strongest within-family pairing (42%), indicating circuits reflect both task demands and model-specific factors. (ii) Component reuse exceeds connection reuse. Models converge on shared functional vocabularies (which components matter) while diverging in routing structure. (iii) Spectral alignment despite topological divergence. Small spectral distances coexist with large edge differences, showing models realise similar information–flow geometry via distinct connectivity. (iv) Scale-dependent refinement within families. Spectral distance decreases monotonically across Pythia scales; cross-family increase reflects architectural differences.
6 Conclusion
Our results characterise semantic-role circuits along three dimensions. RQ1 shows that role information localises into compact, role-specific subgraphs. RQ2 indicates gradual emergence: functional importance can appear early while structural properties evolve, with role-dependent faithfulness. RQ3 reveals partial cross-scale conservation, with components reused across models but connected through different routing schemes. Mechanistically, the absence of sharp transitions suggests that circuits are shaped by sustained optimisation rather than being “switched on” once competence appears. From a linguistic standpoint, these circuits approximate role–filler bindings, central to predicate–argument semantics (who did what to whom, when, with what), implying partially modular causal mechanisms rather than purely diffuse heuristics. COMPASS extends MI temporally (when circuits appear and engage), cross-scale (how they persist across sizes/families), and semantically (predicate–argument relations beyond lexical cues). The presence of compact, causally functional circuits and their partial transfer across models suggests these mechanisms reflect underlying task structure rather than mere memorising surface co-occurrences. For safety and alignment, localising such mechanisms may enable targeted interventions on specific subgraphs, complementing training-time approaches that promote compositionality (aljaafari-etal-2025-carma). Heterogeneous emergence timelines further motivate investigating curricula or early-stopping strategies that prioritise stabilising particular semantic capabilities.
Limitations
@@ 6 Conclusion
This study focuses on English and a subset of roles; whether similar circuits arise in typologically diverse languages or richer role inventories remains unknown. Wide change-point confidence intervals reflect smooth structural trajectories, but more sensitive emergence metrics may yield finer resolution. We analyse decoder-only architectures; encoder–decoder models may exhibit different patterns. Future work should extend to multilingual and multimodal models, study interactions between semantic-role circuits and syntax or coreference mechanisms, and test whether analogous patterns appear for other semantic abstractions such as quantification, modality, or negation.
Ethical considerations
This work aims to improve the interpretability of semantic processing in LLMs through circuit discovery methods. While understanding internal mechanisms benefits safety research and model development, we acknowledge potential risks in localising computational pathways that could be exploited for targeted interventions. Our analysis focuses on controlled semantic tasks with comprehensive validation across training and scales to ensure responsible development and transparent reporting of findings.
@@ B.1 Task: Role-Cross Next-Token Prediction
$\begin{split}\text{Accuracy}=\mathds{1}[y^{(r)}=\operatorname*{arg\,max}_{v}P_{\theta}(v\mid x^{(r)})]\\ \land\mathds{1}[y^{(s)}=\operatorname*{arg\,max}_{v}P_{\theta}(v\mid x^{(s)})]\end{split}$ (2)
@@ B.4 Filtering Procedure
We filter generated pairs to retain only examples where the model predicts the role-appropriate target in both contexts. This ensures circuits are functionally active: the model successfully performs role-specific binding in both clean and corrupted contexts. Examples where either prediction is incorrect are discarded, as they would not reflect active role processing.
@@ B.5 Dataset Statistics
All data is in English and after filtering with models, we obtain:
@@
  • Dual prediction accuracy: 100% (by construction, post-filtering: model predicts the correct target in both clean and corrupted contexts)
  • Cross-role coverage: Each role has at least ${\geq}450$ examples
Table 5 provides a per-role and model breakdown.
Examples per Role Pythia-14M Pythia-410M Pythia-1B LLaMA-1B
Goal 845 895 1052 491
Location 904 673 975 815
Source 667 1206 505 612
Path 802 894 707 959
Instrument 459 1460 $12$ 1098
Beneficiary 773 837 621 1502
Topic 1120 1323 491 1179
Time 557 700 465 712
@@
Appendix C Computation and parameters specifications
C.1 Hyperparameter Selection
Attribution (EAP-IG).
Sparsity and Localization (RQ1).
Emergence Dynamics (RQ2).
Cross-Scale Comparison (RQ3).
Data Filtering.
C.2 Computational Constraints.
C.3 Software Specification.
C.4 Models Specifications.
All models were obtained from Hugging Face wolf2019huggingface and used under their respective intended use, following their respective licenses: Llama 3.2 (Meta Llama 3 Community), Pythia Models (Apache license 2.0), we summarise their key characteristics in Table 6. Pythia models were mainly pre-trained on English data. LLama, on the other hands, has additional multilingual capabilities, including, English, German, French, Italian, Portuguese, Hindi, Spanish, and Thai. The models employ the following tokenisation approaches: Llama 3.2 uses SentencePiece-based BPE, combining 100K tokens from Tiktoken3 with 28K additional tokens to enhance multilingual performance, while Pythia employs GPT-NeoX.
Model Parameters Layers Dmodel Heads Activation
Pythia 14M 1.2M 6 128 4 gelu
Pythia 410M 302M 24 2048 16 gelu
Pythia 1B 805M 16 2048 8 gelu
LLaMA3.2 1B 1.1B 16 2048 32 SiLU
Table 6: Summary of model architectures. Parameters is the total number of trainable parameters; Layers is number of transformer layers; Dmodel: size of word embeddings and hidden states; Heads: number of self-attention heads; and Activation: activation function used in feedforward layers.
Appendix D EAP-IG formalisation
@@
D.1 Transformers as Acyclic Graphs
@@ D.2 EAP-IG Scoring Procedure
    Interpolate inputs: Define the straight-line interpolation $\mathbf{x}(\alpha)=\mathbf{x}^{(r)}+\alpha(\mathbf{x}^{(c)}-\mathbf{x}^{(r)})$ for $\alpha\in[0,1]$.
@@
Intuitively, $S^{\mathrm{IG}}_{u\to v}$ approximates the first-order change in the loss if the contribution of $u$ to $v$’s input were replaced by its corrupted counterpart, averaged along the naturalistic path from corrupt to clean inputs. By integrating over $\alpha$, the estimate reduces sensitivity to local saturation at the clean point and captures non-linear response accumulated along the path. We selected $m=5$, as we saw diminishing returns on higher values, similar to the results in (hanna2024have).
D.3 Score Normalisation and Aggregation.
@@
D.4 Circuits Evaluation metrics
@@
Bootstrap CI for Change-point We estimate a split point $t$ in $(x\!=\!\text{step},\,y\!=\!\text{faithfulness})$ by minimising the summed residuals of two OLS lines on $x_{1},\dots,x_{t}]$ and $x_{t+1},\dots,x_{n}]$ with a minimum segment length of 3. For a 95% CI, we resample pairs $(x,y)$ with replacement, sort by $x$, re-fit the split, map the bootstrap split $x_{t}^{\star}$ back to the original grid, and take the 2.5/97.5 percentiles over 1,000 replicates. We report the point estimate $\hat{t}_{c}$ and the percentile CI over steps.
@@ Appendix E Metric Definitions and Interpretation
Top-$K$ node-mass proportion.
@@
Range $[0,1]$; higher $\Rightarrow$ stronger sparsity. We report $K\in\{5,10,20\}$.
Top-$P$ coverage.
Minimal $K$ such that $\mathrm{TopK}(K)\geq P$, for $P\in\{0.80,0.90,0.95\}$. Lower $K$ indicates higher sparsity.
Gini coefficient (node mass).
Standard Gini on nonnegative masses; range $[0,1]$ ($1=$ all mass on one node). Primary comparator for sparse localisation (RQ1).
@@
Metric Range High Low
Top-$K$ mass $[0,1]$ concentrated circuit (RQ1) diffuse attribution
Top-$P$ coverage ($K$) $\mathbb{N}$ few nodes capture $P$ (sparse) many nodes needed
Gini (mass) $[0,1]$ strong sparsity (RQ1) uniform mass
Density $[0,1]$ saturated links (post-$\hat{t}_{c}$) sparse links
Reciprocity $[0,1]$ feedback motifs feed-forward routing
Avg out-degree $\mathbb{R}_{\geq 0}$ broad fan-out narrow fan-out
Avg weighted out-degree $\mathbb{R}_{\geq 0}$ strong influence spread weak influence
Edge-type mix simplex routing vs. residual balance
Bridges $\mathbb{N}$ bottlenecks (ablation targets) redundancy
Layer span $\mathbb{N}$ deeper integration shallow circuit
Avg betweenness $[0,1]$ (norm.) coordinator hubs flat routing
Top-$K$ Jaccard (step) $[0,1]$ persistent circuit unstable set
Cross-model Jaccard $[0,1]$ architectural consistency family/scale drift
Spectral distance $d_{\text{spec}}$ $\mathbb{R}_{\geq 0}$ similar flow geometry divergent geometry
Table 7: Interpretation guide for graph metrics, with expected high and low values meaning.
@@
Nodes, Edges.
$|V|$ and $|E|$ of the in-circuit graph.
Density.
We use Network’s directed density:
$\begin{split}\mathrm{density}(G)=\frac{|E|}{|V|(|V|-1)}\in[0,1],\end{split}$ (10)
@@
Reciprocity.
Fraction of directed edges participating in reciprocated pairs:
@@
Average out-degree / weighted out-degree.
@@
Edge-type fractions.
@@
Bridges (undirected projection).
Count edges whose removal disconnects the undirected projection of $G$ (structural bottlenecks).
Layer span.
@@
Average betweenness centrality.
$\overline{C_{B}}=\tfrac{1}{|V|}\sum_{v}C_{B}(v)$ on the directed graph (normalized; may use sampling for efficiency).
@@
Detectability $t_{\text{det}}$ (optional).
First step where faithfulness exceeds an early-phase baseline plus $2\sigma$ (baseline/variance from the first $2$ checkpoints), persisting for $\geq 2$ checkpoints.
Indispensability $t_{\text{ind}}$.
Earliest step where ablating the discovered circuit yields a statistically significant performance drop that persists for at least $2$ subsequent checkpoints.
Change-point $\hat{t}_{c}$ (with bootstrap CI).
Two-segment least-squares on $y_{t}\in\{\text{faithfulness},\mathrm{TopK}(K)\}$ with a minimum segment length of $3$. We report $\hat{t}_{c}$ with a nonparametric bootstrap $95\%$ CI.
Consolidation $t_{\text{cons}}$.
@@
Node/edge overlap across models.
Mean Jaccard of Top-$K$ node/edge sets between two models, averaged over common checkpoints. (Defaults: K=30K{=}30.)
Spectral similarity.
For symmetrised, weighted Laplacians ($w_{uv}=|s(u\!\leftrightarrow\!v)|$),
@@
Within-model role overlap.
Jaccard of Top-$K$ nodes between roles at fixed steps (specialisation vs. shared scaffolding).
@@
1:Model checkpoints $\{\theta_{t}\}_{t=0}^{T}$, role-cross dataset $D$, role $r\in\mathcal{R}$, top-$K$ threshold
2:Circuit $\mathcal{C}_{t}^{(r)}$ for each $t$; emergence times $(t_{\text{ind}},\hat{t}_{c},t_{\text{cons}})$
3:Phase 1: Causal Localisation (EAP-IG)
4:for each checkpoint $t=0,\ldots,T$ do
5:  Compute CNP scores $\{\Delta_{\theta_{t}}(y;r,s)\}$ for all $(x^{(r)},x^{(s)})\in D$
6:  Run EAP-IG to obtain edge attributions $\{S_{u\to v}^{\text{IG}}\}_{e\in\mathcal{E}}$ (Appendix D)
7:  Normalise: $\tilde{S}_{u\to v}^{\text{IG}}\leftarrow S_{u\to v}^{\text{IG}}/\sum_{e}|S_{e}^{\text{IG}}|$
8:  Extract circuit: $\mathcal{C}_{t}^{(r)}\leftarrow\text{TopK}(\{|\tilde{S}_{u\to v}^{\text{IG}}|\},K)$
9:end for
10:Phase 2: Temporal Monitoring
11:for each checkpoint $t=0,\ldots,T$ do
12:  Compute faithfulness $F_{t}$ via ablation (See Sec. 3.2)
13:  Compute stability $S_{t}$ via Jaccard (See Sec. 3.2)
14:  Compute structural metrics: Top-$K$ node mass, Gini coefficient
15:end for
16:Phase 3: Emergence Detection
17:Detect indispensability: $t_{\text{ind}}\leftarrow\min\{t:M_{t}(\mathcal{E})-M_{t}(\mathcal{C}_{t})>\epsilon\text{ for }\geq 2\text{ steps}\}$
18:Estimate functional transition: $\hat{t}_{c}\leftarrow\operatorname*{arg\,max}_{t}R^{2}(\text{PiecewiseLinear}(\{F_{t}\}))$ with bootstrap CIs
19:Detect consolidation: $t_{\text{cons}}\leftarrow\min\{t:S_{t}\geq 0.6\text{ for }\geq 2\text{ steps}\}$
20:return $\{\mathcal{C}_{t}^{(r)}\}_{t=0}^{T}$, $(t_{\text{ind}},\hat{t}_{c},t_{\text{cons}})$